THE ORIGIN OF ANOMALOUS ELECTRONIC CIRCULAR DICHROISM SPECTRA OF [RuPt_{2}(tppz)_{2}Cl_{2}]^{4+} IN ACETONITRILE

\underline{H.-G. YU}^{a}, Department of Chemistry, Brookhaven National Laboratory, Upton, NY 11973-5000, USA.

The [RuPt_{2}(tppz)_{2}Cl_{2}]^{4+} (tppz=2,3,5,6-tetra(2-pyridyl)pyrazine) is a potential material for water photo-oxidation to produce oxygen molecules. Recent experiments found that it has anomalous electronic circular dichroism (ECD) spectra in acetonitrile. In order to explain the ECD spectra, we have carried out a detailed study using a hybrid density functional theory (DFT), together with the Stuttgart/Dresden effective core potentials (MWB) for the metal and P atoms. The solvation effects in acetonitrile were taken into account in terms of the conductor polarizable continuum model (C-PCM) with the universal force field (UFF) approach. The UV-vis spectra of the complexes were calculated using the timedependent DFT (TDDFT) method on the optimized geometry of individual system. In this talk, we will discuss the DFT/TDDFT calculations and propose a mechanism for the abnormal ECD spectra.

aThis work was performed at the Brookhaven National Laboratory under Contract No. DE-AC02-98CH10886 with the U.S.

Department of Energy and supported by its Division of Chemical Sciences, Office of Basic Energy Sciences, and also used the resource at NERSC.